OpenStudy (anonymous):

1) y=3x^2+2x x=2 2) y=3x^2+2x x=-4

OpenStudy (bibby):

like before, follow PEDMAS or PEMDAS. (left to right) For 1), plug in 2 so that you get:\[\huge 3(x^2)+2x = 3(2^2)+2(2)\]

OpenStudy (anonymous):

1) 12+4=16

OpenStudy (bibby):

Correct. And 2?

OpenStudy (anonymous):

Is it right :) or :(

OpenStudy (bibby):

:)

OpenStudy (anonymous):

2) 40 right ?

OpenStudy (bibby):

\[\huge 3x^2+2x = 3(-4)^2 + 2 (-4) = (3*16) -8\] recall that a negative number squared is positive and a positive times a negative is a negative

OpenStudy (anonymous):

so is it -40

OpenStudy (bibby):

oh yeah. good job

OpenStudy (anonymous):

Thnx so much

OpenStudy (bibby):

np np

OpenStudy (anonymous):

aww :)